OpenStudy (anonymous):

∫ dt/((e^​t)+1) How do work it out? The answer is: ln| [√(e^t+1) - 1] / [√(e^t+1) + 1] | . Thanks!

OpenStudy (anonymous):

U-Substitution. Set u=e^t, then du=(e^t)dt=udt.

OpenStudy (anonymous):

sorry, but what do you mean by udt?

OpenStudy (anonymous):

Alright, let me be clearer. ∫dt/(e^t+1); problem u=e^t; substitution du=(e^t)dt; evaluating du dt=du/e^t; evaluating dt ∫du/((e^t)(u+1)); sub in u=e^t ∫du/(u(u+1)); remember e^t=u 1=A(u+1)+Bu=Au+A+Bu=(A+B)u+A; partial fractions Can you solve from here?

OpenStudy (anonymous):

ah okay i understand now--thanks!